OpenStudy (anonymous):

let f(x)= 1-2x^2; g(x)=x+1 compute (g of f)(x

OpenStudy (campbell_st):

then its g(1-2x^2) = (1 -2x^2) + 1 = 2 - 2x^2

OpenStudy (anonymous):

now what do i do if I replace x with (-1)

OpenStudy (anonymous):

do I take 2-2(-1)=0

OpenStudy (campbell_st):

just sub x = -1

OpenStudy (anonymous):

thanks